OpenStudy (anonymous):

i dont understand why sometimes you have to use a brachett [] and sometimes ().... how do i know which one to use?

OpenStudy (anonymous):

it's only when you have to put two brackets in a expression, and you want to make it clearer. there's no specific reason, (i think) :P

OpenStudy (amistre64):

brackets group items that beling to a given ...well, group

OpenStudy (amistre64):

you read brackets from the inside out

OpenStudy (amistre64):

if your discussing interval notation tho...

OpenStudy (amistre64):

think of [] as an equals sign connected by a line

myininaya (myininaya):

when doing grouping in algebraic expressions () [] these are them same it doesn't matter which ones you use ------------ () [] these have different notation when talking about sets

OpenStudy (amistre64):

when the interval is EQUAL to, use [ ], otherwise use ( )

OpenStudy (anonymous):

i bet the question is about solving something like \[3<x\leq 5\] solution is \[(3,5]\] closed for less than or equal to, open for strict inequality. that is ( or ) if it is < or > and [ or ] if it is \[\leq\] or \[\geq\]

OpenStudy (anonymous):

so if i dont have the equal to sign then i dont use [] right?

OpenStudy (anonymous):

exactly..

OpenStudy (anonymous):

or to say it positively, if you have equal sign use [ and if you don't use (

OpenStudy (amistre64):

if its equal to, use [ ], otherwise dont

OpenStudy (anonymous):

okay thanks again this explains alot